OpenStudy (anonymous):

how would i solve \[9^{2x+1}=81(27^x)\]

OpenStudy (anonymous):

convert all to base 3 hence we have 3^2(2x+1)=(3^4)(3^(3x)) on multiplication since power gets added we have3^(4x+2)=3^(4+3x) thus 4x+2=4+3x or 4x-3x=4-2 or x=2

OpenStudy (anonymous):

so this would apply to all equations with numbers that have a common multiple then ?

OpenStudy (anonymous):

i would have to convert all the base with the common multiple right?

OpenStudy (anonymous):

ok thanks!